512 g of an unknown metal at a temperature of 15oC is dropped into a 100 g aluminum container holding 325 g of water at 98oC. A short time later, the container of water and metal stabilizes at a new temperature of 78oC. Identify the metal


1
Expert's answer
2021-10-06T15:53:52-0400

Heat gained by metal=heat lost by aluminium+water

"0.512 \times c \times (78-15)=0.1 \times 910 \times (98-78)+0.325\times 4192 \times (98-78) \\

c=900\; J/kg K"

The metal is probably Aluminium.